# Mixed-Integer Programming STALC Stratified Topological Autonomy for Long-Range Coordination

### v1

@@ VI-D3 Start Locations
We add constraints for the start locations of each robot. For each start location $s\in S$, with $n_{s}$ robots at that location, we add the following constraint.
$p_{s,1}=n_{s}$ (28)
VI-D4 Goal Locations
@@
For each goal location, or destination, $d\in D$, we set the minimum number of robots required at that location, $n_{d}$ robots, through the following constraint.
$p_{d,n_{T}}\geq n_{d}$ (29)
VI-D5 Maximum Robots
@@
$\sum_{l\in L}p_{l,t}=n_{A}$ (30)
VI-D6 Sequential Flow
We add a sequential flow constraint to ensure that movement across the graph is restricted to the structure of the graph. In particular, for each node, the number of robots in the node and flowing into the node must be equal to the number of robots in the node and flowing out of the node in the next time step. Thus, for each time $t\in[2,n_{T}]$ and node $v_{j}$, we add the following constraint.
$\sum_{\begin{subarray}{c}l_{ij}=(v_{i},v_{j})\in L\end{subarray}}p_{l_{ij},t-1}=\sum_{\begin{subarray}{c}l_{ji}=(v_{j},v_{i})\in L\end{subarray}}p_{l_{ji},t}$ (31)
The first sum considers all locations of the form $l_{ij}=(v_{i},v_{j})$ and the second sum considers all locations of the form $l_{ji}=(v_{j},v_{i})$ for a fixed node $v_{j}$. Both sets of locations include $l_{jj}$ since all nodes have self-loops. This constraint assumes a robot would never stay on an edge (due to the cost). This is enforced by the cost function as long as the cost of all edges is greater than zero and overwatch is never provided from an edge. This constraint allows robots to move from one edge to the next without stopping at the node.
VI-D7 Negative Edge Weights
We add a constraint to ensure that edge weights do not become negative due to the reductions from overwatch opportunities. For each edge and time step, we add the following constraint.
$\breve{C}_{W_{e,t}}+\sum_{{o}=(v_{i},e)\in{O}}\breve{C}_{\Omega_{{o},t}}\geq\phi_{e,t}$ (32)
When the edge is used ($\phi_{e,t}=1$), this constraint ensures that the minimum cost associated with the edge is $1$, and otherwise 00. The cost is strict to this bound since we are minimizing costs. This constraint is only necessary for edges that can be monitored through overwatch and can be omitted for the other edges if there are checks in place for the relative values of the edge weight to teaming benefits.
VI-E MIP Optimization Problem
Combining our objective function and constraints from the previous sections, our overall MIP optimization problem is expressed in Table III.
@@ VI-E MIP Optimization Problem
$\min~{}\sum\limits_{t=1}^{n_{T}}\bigg{(}C_{T_{t}}+\sum\limits_{e\in E}\breve{C}_{W_{e,t}}+\sum\limits_{{o}\in{O}}\breve{C}_{\Omega_{{o},t}}\bigg{)}$ subject to (25)
Cost Constraints $\breve{C}_{W_{e,t}}\geq-m_{e}p_{e,t}+(w_{e}+m_{e}a_{e})\phi_{e,t},$ $\forall e,t$ (16)
$\breve{C}_{W_{e,t}}\geq-r_{e}p_{e,t}+(w_{e}+r_{e}a_{e})\phi_{e,t},$ $\forall e,t$ (17)
@@
$\breve{C}_{W_{e,t}}+\sum\limits_{{o}=(v_{i},e)\in{O}}\breve{C}_{\Omega_{{o},t}}\geq\phi_{e,t}$ $\forall e,t$ (32)
By employing our key formulation innovations from Sec. VI-B, we can use mixed-integer linear programming to solve our problem since we have a linear objective function with linear constraints.
We solve the optimization program shown in Table III using the Gurobi optimizer [39].
VI-F MIP Problem Formulation Considerations
We formulate our optimization problem such that edges have cost and nodes do not. In particular, we assure that there is not a benefit to staying on an edge, such that any movement on edge (and associated cost accrual) is necessary to achieve the overall objective. However, remaining at a node could provide a benefit, such as to enable performing overwatch of other traversing team members. This assumption is built into the problem’s formulation.
@@ VI-F MIP Problem Formulation Considerations
We assume a graph without negative cycles. Overwatch for one edge can come from multiple nodes, but the weight of the edge (resulting from the cost of traversing, benefit of overwatch, and cost reductions from vulnerability/teaming) cannot reduce to 0 or below. We perform checks on the overwatch reductions to ensure that these conditions are met. Additionally, we add the constraint in equation (32) to bound the edge weights. The formulation proposed herein allows robots at one node to overwatch robots traversing multiple edges. Intuitively, for an overwatch opportunity to be used, the cost reduction from overwatch needs to be more than the cost to get to the overwatch position. Otherwise, the robots will stay together for teaming benefits.
@@
While our solutions are always guaranteed to be optimal, the optimal solution is not guaranteed to be unique. Small problems with simple numbers can result in many equivalent cost solutions and cause computation time to increase. For example, with teaming, without any other factors such as vulnerability and overwatch, if the $r_{e}$ cost reductions are equivalent on all edges and there are multiple routes being taken, more robots will go on the longest route, which is logical from a risk mitigation perspective. If route lengths are equal, multiple optimal solutions will arise. We demonstrate simple simulated scenarios where this is more of a consideration than in our hardware experiments where we utilize our environment segmentation approach.
@@ VII Mid-Level Planner
Our mid-level planner takes the output from the MIP problem, determines between which nodes each robot is intended to traverse and generates minimally visible paths to follow.
VII-A Robot Route Assignment
After generating an optimal plan from our optimization problem in Table III, we use the resulting decision variable $p_{l,t}$ in an assignment routine to determine the routes for each robot through the graph. We use each start location from our set $S$ and then go to locations in the form $(v_{1},v_{2})$. At each time step, we assign a robot to a location with a start point, $v_{1}$, that matches their current end point, $v_{2}$, to maintain movement along the graph. For a node, the start and end points are equal (i.e., $v_{1}=v_{2}$). We do this for each robot until all robots have a route through the graph. We present Algorithm 3 with the details of this assignment process. We define the following functions for our algorithm: NewList() to create a new empty list, RemoveItem($S$) to remove an item from set $S$ (and output that item), and AddItem($m$) to add item $m$ to an existing list. Tracking the number of robots at each location rather than routes for each robot directly necessitates this post-processing step, however the paradigm shift to integer decision variables (as described in Sec. VI-B1) enables a significant reduction in computational complexity by removing the dependence on the number of robots in the size of our decision space.
Algorithm 3 Robot Route Assignment
0:

### v2

@@ V-C3 Start Locations
We add constraints for the start locations of each robot. For each start location $b\in B$, with $n_{b}$ robots at that location, we add the following constraint:
$p_{b,1}=n_{b}.$ (21)
V-C4 Goal Locations
@@
For each goal location, or destination, $d\in D$, we set the minimum number of robots required at that location, $n_{d}$ robots, through the following constraint:
$p_{d,n_{T}}\geq n_{d}.$ (22)
V-C5 Maximum Robots
@@
$\sum_{l\in L}p_{l,t}=n_{A}.$ (23)
V-C6 Sequential Flow
We add a sequential flow constraint to ensure that movement across the graph is restricted to the structure of the graph. In particular, for each node, the number of robots in the node and flowing into the node must be equal to the number of robots in the node and flowing out of the node in the next time step. Thus, for each time $t\in[2,n_{T}]$ and node $v_{j}$, we add the following constraint:
$\sum_{\begin{subarray}{c}l_{ij}=(v_{i},v_{j})\in L\end{subarray}}p_{l_{ij},t-1}=\sum_{\begin{subarray}{c}l_{ji}=(v_{j},v_{i})\in L\end{subarray}}p_{l_{ji},t}.$ (24)
The first sum considers all locations of the form $l_{ij}=(v_{i},v_{j})$ and the second sum considers $l_{ji}=(v_{j},v_{i})$ for a fixed node $v_{j}$. Both sets of locations include $l_{jj}$ since all nodes have self-loops. This constraint allows robots to move from one edge to a node or to another edge without stopping at the node.
V-C7 Negative Edge Weights
We add a constraint to ensure that edge weights do not become negative due to the reductions from overwatch opportunities. For each edge $e$ and time step $t$, we add the following constraint:
$\breve{C}_{\bar{W}_{e,t}}+\sum_{{o}=(v_{i},e)\in{O}}\breve{C}_{\Omega_{{o},t}}\geq\phi_{e,t}.$ (25)
When the edge is used ($\phi_{e,t}=1$), this constraint ensures that the minimum cost associated with the edge is 11. The cost is strict to this bound since we are minimizing costs.
V-D MIP Optimization Problem
Combining our objective function and constraints from the previous sections, our overall MIP optimization problem is expressed in Table III. By employing our key formulation innovations from Sec. V-A, we can use mixed-integer linear programming (MILP) to solve our problem since we have a linear objective function with linear constraints. We use the Gurobi optimizer [60] for solving this problem to optimality.
@@ V-D MIP Optimization Problem
$\min~\sum\limits_{t=1}^{n_{T}}\bigg(C_{T_{t}}+\sum\limits_{e\in E}\breve{C}_{\bar{W}_{e,t}}+\sum\limits_{{o}\in{O}}\breve{C}_{\Omega_{{o},t}}\bigg)$ subject to (18)
Cost Constraints $\breve{C}_{\bar{W}_{e,t}}\geq-m_{e}p_{e,t}+(\bar{w}_{e}+m_{e}a_{e})\phi_{e,t},$ $\forall e,t$ (9)
$\breve{C}_{\bar{W}_{e,t}}\geq-r_{e}p_{e,t}+(\bar{w}_{e}+r_{e}a_{e})\phi_{e,t},$ $\forall e,t$ (10)
@@
$\breve{C}_{\bar{W}_{e,t}}+\sum\limits_{{o}=(v_{i},e)\in{O}}\breve{C}_{\Omega_{{o},t}}\geq\phi_{e,t}$, $\forall e,t$. (25)
V-E MIP Problem Formulation Considerations
@@
We formulate our optimization problem assuming edges have cost and nodes do not, such that there is not a benefit to staying on an edge. Additionally, we require that costs for not meeting a minimum number of robots on an edge are set to be greater than reductions for agents after that number (i.e., $m_{e}\geq r_{e}$), as noted in Sec. V-B1. We assume a graph without negative cycles. Overwatch for one edge can come from multiple nodes and robots at one node can overwatch robots traversing multiple edges, but the weight of the edge (resulting from the cost of traversing, benefit of overwatch, and cost reductions from vulnerability/teaming) cannot reduce to 0 or below. We enforce this property with the constraint in (25). Intuitively, for an overwatch opportunity to be used, the cost reduction from overwatch needs to be more than the cost to get to the overwatch position.
@@
We solve problems with our MIP formulation to optimality. However, while our MIP solutions are always guaranteed to be optimal, the optimal solution is not guaranteed to be unique. Small problems with simple numbers can result in many equivalent cost solutions and cause computation time to increase. We demonstrate simple simulated scenarios with simple numbers where this is more of a consideration than in our hardware experiments where we utilize our graph generation approach.
VI Graph Generation
Constructing a graph that reflects the topology and path costs of the underlying metric space is essential for unifying the graph-based coordination problem in (2) and the local multi-robot motion planning described in III-B. Because the graph is inherently connected to the problem domain, we focus specifically on constructing a graph for the MCRP.
VI-A Visibility Map
Since the MCRP aims to minimize risk due visibility, the graph is generated by first constructing a visibility cost map. We accomplish this by adapting the notion of a viewshed[61] to the context of a distribution over observer positions. The viewshed of an area is a binary mask marking the regions of a Digital Elevation Model (DEM) that are visible from a given observer position. Our goal is to estimate the probability that each point in the environment can be seen by the observer.
Line-of-sight analyses, such as viewsheds, are commonly used in terrain analysis for tactical decision making (e.g., in [47, 48] for determining cover and concealment). Uncertainty in terrain measurements and observer locations can significantly impact the accuracy of viewshed calculations. While we do not address terrain uncertainty in this work, we formulate a probabilistic notion of a viewshed that incorporates uncertainty in the observer position.
[…]
We refer to the map containing the probability of being seen at each point as the visibility map. An example of a visibility map is depicted in Fig. 4 with a single observer. The lighter regions indicate higher visibility.
Equation 26 does not consider the distance from the observer when computing the probability of detection. In practice, however, distant objects appear less visible. We model the visibility of an object as proportional to its apparent angular size, $\theta$, which scales inversely with distance, $d$, from the observer (i.e., $\theta\sim\frac{1}{d}$). The distance is determined using the distance transform to compute the Euclidean distance from each map cell to the nearest observer distribution. For distributions with infinite support (e.g., Gaussian), we must choose a “boundary” for the distribution in order to compute the distance transform. In this work, we use the 2-$\sigma$ ellipse. Detection probability is then scaled by the ratio of the nearest enemy distance to the maximum effective distance $d_{\textrm{max}}$. $d_{\textrm{max}}$ is chosen based on the observer characteristics (e.g., when the robot’s angular size falls below a threshold or spans fewer than a set number of pixels in a camera image). We define the distance-weighted visibility map, $\mathcal{P}_{\textrm{vis}}(x,y)$ for a position $(x,y)$, where distance to the nearest observer is $d(x,y)$.
[…]
Algorithm 1 Construct Topological Graph
1:
$\mathcal{D}$ Digital Elevation Model
$C_{m}$ Obstacle map
$\mathcal{O}$ Observer distribution
NN Number of samples from observer distribution
$\nu$ Visibility threshold
$\Xi_{\textrm{min}}$ Minimum cover region size
$\Xi_{\textrm{max}}$ Maximum cover region size
$\lambda_{p}$ Visibility cost weight
2:
$V$ Graph nodes (cover region centroids)
$E$ Graph edges
$W_{\textrm{vis}}$ Visibility weight adjacency matrix
$\mathcal{W}_{\textrm{ow}}$ Set of overwatch weight matrices, $W_{\textrm{ow}}(v)$
[…]
12:for all $\{(v_{j},v_{k}):\mathcal{T}\}\in\mathcal{T}_{\textrm{feas}}$ do
[…]
15://// Overwatch Opportunities
16:for $v\in V$ do
[…]
20:for all $\{(v_{j},v_{k}):\mathcal{T}\}\in\mathcal{T}_{\textrm{feas}}$ do
[…]
22:return $V,E,W_{\textrm{vis}},\mathcal{W}_{\textrm{ow}}$
Algorithm 2 ComputePaths
1:
$\mathcal{P}_{\textrm{vis}}$ Visibility map
$C_{m}$ Obstacle map
$\mathcal{R}$ Cover regions
$V$ Graph nodes
$\lambda_{p}$ Visibility cost weight
2:
$\mathcal{T}_{\textrm{feas}}$ Map: $(v_{j},v_{k})\mapsto\mathcal{T}$
[…]
8:for all $v_{j}\in V$ do
9:for all $v_{k}\in V$ do
10:  if $v_{j}=v_{k}$ then
11:   continue
12:  $\mathcal{T}\leftarrow$ ComputeOnePath($C_{m}$, $\mathcal{P}_{\textrm{vis}}$, $v_{j}$, $v_{k}$, $\lambda_{p}$)
13:  if PathIsRedundant($\mathcal{R}$, $\mathcal{T}$, $v_{j}$, $v_{k}$) then
14:   $path\_to[v_{k}]\leftarrow$ MinCost($\mathcal{T}$, $path\_to[v_{k}]$)
15:   $path\_from[v_{j}]\leftarrow$ MinCost($\mathcal{T}$, $path\_from[v_{j}]$)
16:  else
17:   $\mathcal{T}_{\textrm{feas}}[(v_{j},v_{k})]\leftarrow\mathcal{T}$
18:   $is\_start[v_{j}]\leftarrow$ true
19:   $is\_goal[v_{k}]\leftarrow$ true
[…]
21:            path_frompath\_from, path_topath\_to)
22:return $\mathcal{T}_{\textrm{feas}}$
VI-B Constructing a Topological Graph
Using Def. III.1 for a DTG, the graph $G=(V,E,W_{E})$ is composed of nodes $V$, edges $E$, and weights $W_{E}$ that are a function of the state of the robots. For the MCRP problem, we decompose the edge weights $W_{E}$ into two components: visibility edge weights $W_{\textrm{vis}}$ and overwatch edge weights $\mathcal{W}_{\textrm{ow}}$. As used in Sec. V-B, $\bar{w}_{e}\in W_{\textrm{vis}}$ and $\omega_{o}\in\mathcal{W}_{\textrm{ow}}$. We define $W_{\textrm{vis}}$ to be a weighted adjacency matrix and $\mathcal{W}_{\textrm{ow}}$ is a set of weighted adjacency matrices for each node, $\mathcal{W}_{\textrm{ow}}=\{W_{\textrm{ow}}(v):v\in V\}$, where element $W_{\textrm{ow}}(v)(j,k)$ denotes the overwatch that a team of robots positioned at node $v$ can provide for a team traversing edge $(j,k)$. Algorithm 1 presents our procedure for constructing the components of the topological graph. This algorithm references the function ComputePaths defined in Algorithm 2. The following subsections describe the key steps (represented in both algorithms as functions) in more detail.
VI-B1 ComputeVisMap
The visibility map computation uses an efficient viewshed algorithm [61] to generate NN viewshed maps sampled from the observer distribution $\mathcal{O}$. Equations (26) and (27) are used to compute the probability of detection for each cell in the map. These calculations form our visibility map $\mathcal{P}_{\textrm{vis}}$.
VI-B2 GetCoverMask
The cover calculation creates a mask of all cells which have low visibility and are traversable.
[…]
VI-B3 GetCoverRegions
We compute cover regions, $\mathcal{R}=\{r\in\texttt{ConnComp}(\mathcal{M}_{c}):\texttt{area}(r)>\Xi_{\textrm{min}}\}$, by finding all of the discrete, connected components of the cover mask whose areas exceed a threshold value, $\Xi_{\textrm{min}}$. For this, we use functions labeled ConnComp and area, respectively.
(a) Cover regions on visibility map
(b) Pruned A paths between cover regions
Figure 5: Regions of cover and A paths overlaid on satellite imagery of the environment: cover regions are outlined in red and node numbers are located within the regions near the centroid of each region.
VI-B4 SplitRegions
For some maps and observer positions, the cover regions can be large and nonconvex. Consequently, we enforce a maximum size, $\Xi_{\textrm{max}}$, using the algorithm from [62] to find the minimum-length cut that divides one region into two smaller ones. The result of this cut is 2 regions, where one region is of size $\Xi_{\textrm{max}}$ and the other is the remainder. We perform these cuts repeatedly until no region is larger than $\Xi_{\textrm{max}}$. Other strategies for splitting regions could be used; empirically, we find that iterative minimum length cuts tend to produce reasonably compact regions.
VI-B5 PlaceNodes
For each region, we place a node at the closest point in the region to its centroid. Since the regions can be nonconvex, the centroid is not guaranteed to be contained in the region. The following defines the mapping $M_{v}$ between regions and nodes: $V=\{\,\operatorname*{arg\,min}_{p\in r}(\|p-\texttt{centroid}(r)\|_{2}):r\in\mathcal{R}\,\}.$ Here, the function centroid finds the centroid of the region. In Fig. 5(a), we overlay outlines of the cover regions in an example environment.
VI-B6 ComputePaths
We present Algorithm 2 for computing the feasible paths, $\mathcal{T}_{\textrm{feas}}$. $\mathcal{T}_{\textrm{feas}}$ maps from nodes $(v_{j},v_{k})$ to a path $\mathcal{T}$ between those nodes. Empty maps are initialized with the function EmptyMap. The resulting paths are composed of spatial locations: $\mathcal{T}=\{\tau_{1},\dots,\tau_{n}\}$, $\tau_{i}=[x_{i},y_{i}]^{T}$.
VI-B7 ComputeOnePath
We use an optimal path planning algorithm to compute paths that minimize visibility between each pair of nodes in the graph. In this work, we use a standard A search. We choose A for its optimality and simplicity, but in principle any optimal path planner capable of planning over costmaps could be used. We define the heuristic cost, $h(\tau)$, as the Euclidean distance from the current position to the goal. The function $g(\tau)$ denotes the cost to follow the optimal path from the start position to $\tau$. The successor cost, $c(\tau,\tau^{\prime})$, is the cost to go from a cell in the map $\tau$ to adjacent cells $\tau^{\prime}$. Finally, $n(\tau)$ is the visibility cost of traversing a single cell, which we will define in (33).
[…]
$g(\tau^{\prime})=g(\tau)+c(\tau,\tau^{\prime}).$ (30)
In (29), $\lambda_{p}$ is a scaling factor to weight the contribution of the visibility cost. Since the true cost-to-go is equal to the heuristic cost plus a non-negative visibility term, the heuristic $h(\tau)$ is trivially admissible.
VI-B8 PathIsRedundant
When computing the set of graph edges, we initially consider the topological graph to be fully-connected. However, many of these edges will overlap due to minimizing visibility or the routing of the environment, so we take an additional step to eliminate redundant edges. We define an edge $e=(v_{j},v_{k})$ as redundant if its associated path $\mathcal{T}_{v_{j}v_{k}}$ passes through a region other than the two regions it is intended to connect (i.e., $\mathcal{T}_{v_{j}v_{k}}\cap\big(\bigcup_{r\in\mathcal{R}\setminus\{r_{j},r_{k}\}}r\big)\neq\varnothing$).
Whenever a redundant edge is removed, we update the minimum cost (with function MinCost) of the pruned paths originating at $r_{s}$ and ending at $r_{g}$ for later use in ReconnectNodes.
VI-B9 ReconnectNodes
In rare circumstances, the pruning step may result in a node which is a sink (i.e., not the start of any path), a source (i.e., not the goal of any path), or fully disconnected. In these cases, we restore the lowest cost edges needed to maintain a directed graph with no sources or sinks. If there is not a viable route to a particular node (e.g., due to obstacles in the environment), that node will remain disconnected and will not be used in the graph planning process. An example of the pruned paths is shown in Fig. 5(b).
VI-B10 PathCost
When traversing a path $\mathcal{T}$ through the environment, we consider the total probability of non-detection, $p_{nd}$, as the product of the probabilities of non-detection at each cell (assuming independence). We use a general $\mathcal{P}$ to represent a visibility map $\mathcal{P}_{\textrm{vis}}$ or overwatch map $\mathcal{P}_{\textrm{ow}}(v)$ (which we will define in Sec. VI-B11).
$p_{nd}=\prod_{\tau\in\mathcal{T}}(1-\mathcal{P}(\tau)).$ (31)
Taking the negative $\log$ of this quantity, we can compute a negative $\log$ non-detection probability by summing the contributions of each cell along the path.
$-\log(p_{nd})=\sum_{\tau\in\mathcal{T}}-\log(1-\mathcal{P}(\tau)).$ (32)
As the probability of detection approaches one, the cost to traverse a cell approaches infinity. In order to avoid infinite values in the optimization problem, we bound the maximum probability of detection at each cell to $1-\epsilon$ (for a small positive $\epsilon$). So the cost to traverse a single cell is the negative-log of the probability of non-detection.
$n(\tau)=-\log(\max(1-\mathcal{P}(\tau),\epsilon)).$ (33)
In the final step of the path cost calculation, we compute the weight for each edge $e=(v_{j},v_{k})$ that remained after the pruning step by considering the optimal path $\mathcal{T}_{v_{j}v_{k}}$.
$W(v_{j},v_{k})=\sum_{\tau\in\mathcal{T}_{v_{j}v_{k}}}n(\tau).$ (34)
VI-B11 Overwatch Opportunities
The overwatch weighted adjacency matrix, $W_{\textrm{ow}}(v)$, represents the benefit of overwatch from node $v$. Computing this matrix is analogous to the visibility map calculation. Instead of sampling from the observer distribution, we sample uniformly from the cover region associated with node $v$ (i.e., $r_{v}$). We define the uniform distribution (with function UniformDistribution) associated with $r_{v}$ as $\mathcal{O}_{\textrm{ow}}(v)$. We apply equations (26) and (27) to produce a visibility map for an observer located at node $v$, which we call an overwatch map, $\mathcal{P}_{\textrm{ow}}(v)$. The overwatch edge weights, $W_{\textrm{ow}}(v)$, are computed using the PathCost function to evaluate the visibility from the node to each edge in the graph. The overwatch edge weight scales with how visible a particular edge is from the overwatch node and the distance along the edge that is visible. Thus, higher edge weights correspond to more advantageous overwatch opportunities. An example overwatch map is shown in Fig. 6.
VI-C Topological Graph Refinement
We graphically depict the information embedded in our topological graph through the schematic representation shown in Fig. 7. The graph includes nodes for each cover region, edges representing the pruned A paths, and overwatch opportunities from the overwatch maps, denoted by arrows from the overwatch node to the edge that can be monitored. Although our formulation considers a directed graph, we construct graphs with both edge directions; thus, our schematics are shown with undirected edges.
When transforming the environment segmentation into this schematic representation, we use heuristics to refine the graph based on the robots’ capabilities and the scenario. In particular, we set a maximum distance that can be traversed and remove edges from $E$ that do not meet this threshold.
We add overwatch opportunities from $\mathcal{W}_{\textrm{ow}}$ to our graph after applying scenario-based heuristics. First, we impose a maximum overwatch distance (e.g., from sensor limits) by checking distances to edge endpoints. We then add a scale factor to the overwatch weights based on the priorities of the scenario (e.g., if overwatch is more or less valuable). Finally, we only incorporate overwatch opportunities when the cost reduction exceeds a prescribed fraction of the edge weight, and we bound larger reductions to ensure traversal always incurs nonzero cost (e.g., overwatch could provide between a 40% and 90% cost reduction). These thresholds and scaling factors are scenario-dependent and we demonstrate different values for these parameters in our analysis.
@@ VII Mid-Level Planner
Our mid-level planner serves to connect the graph planner solutions of (2) to the local multi-robot motion planning problem in III-B by assigning robots to edges and generating robot-specific mid-range plans.
VII-A Robot Route Allocator
After generating an optimal plan from our optimization problem in Table III, we use the resulting decision variable $p_{l,t}$ in an assignment routine to allocate routes for each robot through the graph. We use each start location from our set $B$. At each time step $t$, each robot $i$ will move to states in the form $s_{t}^{i}=(v_{j},v_{k})$. Subsequent valid states are assigned from the Next Edge Action Set $A(s_{t}^{i})$ defined in Def. III.2. All nodes have self loops (i.e., $v_{j}=v_{k}$ for nodes). We present Algorithm 3 with the details of this assignment process. We define the following functions for our algorithm: RemoveItem($B$) to remove an item from set $B$ (and output that item) and AddItem($m$) to add item $m$ to an existing list. Tracking the number of robots at each location rather than routes for each robot directly necessitates this post-processing step, however the paradigm shift to integer decision variables (as described in Sec. V-A1) enables a significant reduction in computational complexity by removing the dependence on the number of robots in the size of our decision space.
Algorithm 3 Robot Route Allocator
1: